ESPERANTOLOGIO

ESPERANTO STUDIES

EES

Kajero 2  2001  Issue No. 2

Redaktas Editor

Christer Kiselman

Redakta komitato Editorial Board

Detlev Blanke José Borges Neto Probal Dasgupta Michel Duc Goninaz Gotoo Hitosi Edmund Grimley Evans Koutny Ilona Erich-Dieter Krause Sergej Kuznecov Jouko Lindstedt Liu Haitao Ingrid Maier Geraldo Mattos Marc van Oostendorp Sergio Pokrovskij Humphrey Tonkin John C. Wells Yamasaki Seikô

Varna

Upsala Universitato Uppsala University

ISSN 1311-3496 (p) ISSN 1404-4709 (e)

Presita en Bulgario de / Printed in Bulgaria by

Color-Print

Varna 2001